OpenStudy (wolfman77):

10th grade help! Math! MEDAL and FAN!!! Emmanuel bought 28 lb of potting soil this week. This amount is 4 lb more than twice the amount of potting soil he bought last week. Answer the following questions to find the number of pounds of soil Emmanuel purchased last week. (a) What is the unknown information? (b) Let p represent the unknown information. Write an equation to model the problem. (c) Which number, 10, 11, or 12, will solve the equation in Part (b)? Show your work. (d) How many pounds of potting soil did Emmanuel buy last week?

OpenStudy (wolfman77):

@Jadeishere @MrNood

OpenStudy (wolfman77):

@mathstudent55

OpenStudy (jadeishere):

So we're told that he bought 28 lbs of soil this week, and that it's 4 more lbs than last weeks amount. What do we not know, from the given information? We're told he bought some this week and last week, what are we not told?

OpenStudy (wolfman77):

Were not told the amount of lb's of soil from last wk.

OpenStudy (jadeishere):

Exactly, so 'p' represents that amount. We need to find out how much was bought last week. So twice the amount + 4 of last weeks soil is equal to 28 Do you know how we'd write that in an equation?

OpenStudy (wolfman77):

28 - 4 = p ?

OpenStudy (wolfman77):

Or p + 4 = 28

OpenStudy (jadeishere):

Not quite, if it's twice the amount of last weeks soil, plus 4 what would we do to make sure it's TWICE the amount + 4?

OpenStudy (wolfman77):

2p + 4 = 28 ?

OpenStudy (jadeishere):

Yes! That's your equation that you have to use :) Great job!

OpenStudy (wolfman77):

Thanks! Next...

OpenStudy (jadeishere):

Kay, now we need to simplify the equation as though it's a regular equation 1) isolate the variable 2p + 4 = 28 2p = 28 - 4 2p = 24 2p / 2 = 24 /2

OpenStudy (jadeishere):

So divide both sides by 2 (what you do to one side, you must do to the other)

OpenStudy (wolfman77):

24 / 2 = 12

OpenStudy (wolfman77):

So, 2p would be 12 x 2 right?

OpenStudy (jadeishere):

Yes, divide BOTH sides by 2 though so 2p divided by 2 = p p=12 :)

OpenStudy (wolfman77):

And for the last part, the answer would be: He bought 24 lbs last wk?

OpenStudy (jadeishere):

No, last week he bought 12

OpenStudy (wolfman77):

Yeah, I meant that.

OpenStudy (wolfman77):

Thanks!!!!!!!!!!!!!!!!!!

OpenStudy (jadeishere):

Yup, great job! Just throwing this out there, please don't feel obligated to say that you're a girl in order to get help.. It's demeaning and sad to see that that's the only way people can get help on here :/ If you ever need help, just tag me and I'll do my best :) ~Jade Amaia Rai

OpenStudy (wolfman77):

I was so tired of everyone calling me a boy, that's why I put tht, but thanks!!